What mass ofPCl

5

P

C

l

5

 will be produced from the given masses of both reactants?

Express your answer to three significant figures and include the appropriate units.

View Available Hint(s)



1
Expert's answer
2021-11-17T04:18:01-0500

Solution

P4 + 10 Cl2 = 4 PCl5

Combination of one mole of P4 with ten moles of chlorine gives four moles of PCl5.

MW(P4) = 123.895048 g mol-1

MW(Cl2) = 70.906 g mol-1

MW(PCl5) = 208.22 g mol-1

Part B

n(P4) = m(P4)/MW(P4) = 26.0 g/123.895048 g mol-1 = 0.210 mol

n(PCl5) = 4 n(P4) = 0.839 mol

Part C

N(Cl2) = m(Cl2)/MW(Cl2) = 59.0 g/70.906 g mol-1 = 0.832 mol

n(Cl5) = 0.4 n(Cl2) = 0.333 mol

Part D

Chlorine is limiting reagent. 0.333 moles of PCl5 will be formed:

m(PCl5) = n(PCl5)*MW(PCl5) = 0.333 mol * 208.22 g mol-1 = 69.3 g